How many ounces of a 17% acid solution and a 56% acid solution are needed to obtain 10 ounces of a 34% acid Mixture?
Answer by jim_thompson5910(35256)

Hint: solve the system for x and y

x+y = 10

0.17x + 0.56y = 10*0.34


where x = # of ounces of the 17% acid solution and y = # of ounces of the 56% acid solution